# Feedback-Free Resource Scheduling: Towards Flexible Multi-BS Cooperation in FD-RAN

### v1

@@ Abstract
Flexible cooperation among base stations (BSs) is critical to improve resource utilization efficiency and meet personalized user demands. However, its practical implementation is hindered by the current radio access network (RAN), which relies on the coupling of uplink and downlink transmissions and channel state information feedback with inherent issues such as overheads and delays. To overcome these limitations, we consider the fully-decoupled RAN (FD-RAN), in which uplink and downlink networks are independent, and feedback-free MIMO transmission is adopted at the physical layer. To further deliver flexible cooperation in FD-RAN, we investigate the practical scheduling process and study feedback-free downlink multi-BS resource scheduling. The problem is considered based on network load conditions. In heavy-load states where it is impossible for all user demands to be met, an optimal greedy algorithm is proposed, maximizing the weighted sum of user demand satisfaction rates. In light-load states where at least one solution exists to satisfy all user demands, an optimal two-stage resource allocation algorithm is designed to further minimize network energy consumption by leveraging the flexibility of cooperation. Extensive simulations validate the superiority of proposed algorithms in performance and running time, and highlight the potential for realizing flexible cooperation in practice.
@@ I Introduction
However, implementing flexible cooperation confronts significant challenges since the architecture of the fifth-generation (5G) radio access network (RAN) still follows the original single cell based design [6]. On the one hand, the coupling between uplink (UL) and downlink (DL) transmissions restricts the flexibility of BS association and resource utilization. For example, in time division duplex (TDD) systems, cooperating BSs are required to maintain consistent UL and DL frame structures, and it is not feasible to independently configure different cooperation sets for UL and DL transmissions. Moreover, the performance gains of cooperation heavily depend on the accuracy and timeliness of channel state information (CSI), which can only be ensured in systems with tightly coupled UL and DL. On the other hand, the overheads of pilots and CSI feedback scale linearly with the number of cooperating BSs and users, reducing the effective data rate, as flexible cooperation requires CSI for arbitrary pair of BS and user. Furthermore, the computational complexity of determining joint transmission parameters for multi-BS cooperation increases exponentially [7]. As a result, practical cooperation is generally constrained to two fixed BSs [8]. Besides, inherent feedback delays in transmission lead to outdated CSI, which will deviate the cooperative resource scheduling decisions from its desired objectives, e.g., meeting specific users’ demands.
@@
However, there still lacks existing research on feedback-free resource scheduling towards flexible multi-BS cooperation. To this end, we investigate the problem of DL cooperation set selection and subcarrier allocation in FD-RAN. The flexibility of resource allocation lies in that users can be associated with any potential DL-BSs operating in appropriate TMs. However, such flexibility also significantly increases the complexity of the problem, posing challenges to achieving optimal solutions. Furthermore, resource allocation can have diverse and even conflicting objectives, including network throughput, fairness, and energy efficiency (EE). Thus, appropriate scheduling objectives are required for various conditions to better utilize resources. Moreover, to align with the practical scheduling process, resource scheduling should meet stringent running time requirements, e.g., at the ms level. Consequently, designing resource allocation algorithms that both realize computational efficiency and scheduling quality in FD-RAN remains a very challenging problem.
To address the above challenges, we propose a feedback-free cooperative resource scheduling framework that can leverage the flexibility to achieve tailored scheduling objectives based on the current network load condition. We also consider resource scheduling from the user’s perspective, focusing on meeting personalized service demands. Specifically, the network is classified into heavy-load and light-load conditions based on whether all user demands can be met through JT of all DL-BSs, which provides the highest data rate per subcarrier. In heavy-load conditions, where user demands cannot be fully met, JT is still employed to maximize demand satisfaction. A greedy subcarrier allocation algorithm considering user service priorities and fairness is proposed and proves optimal. In light-load conditions, where at least one solution exists to satisfy all user demands by JT of all DL-BSs, the focus shifts to finding the most energy-efficient cooperation strategy. An efficient two-stage resource allocation (TSRA) algorithm is proposed, which first employs the feasibility pump algorithm [16] to quickly obtain a feasible solution and then iteratively refines it to optimality using the branch-and-cut method [17]. To conclude, the contributions of this paper are listed as follows:
@@
  • To solve the problems, the resource scheduler is designed with two tailored algorithms. In heavy-load conditions, a greedy subcarrier allocation algorithm is proposed and proves optimal. In light-load conditions, the TSRA algorithm is proposed to enable flexible and energy-efficient cooperation. A feasible solution is quickly obtained using the feasibility pump algorithm and then iteratively refined to achieve optimality through the branch-and-cut method.
@@ II-B Resource Allocation Towards Multi-BS Cooperation
Over the past decade, substantial efforts have been dedicated to resource allocation aimed at enabling multi-BS cooperation across diverse network settings and objectives[24, 25, 26, 27, 28, 29]. [24] proposes a coordinated scheduling, beamforming, and power allocation scheme across multiple BSs in wireless cellular networks. The study primarily aims to mitigate inter-cell interference and enhance overall network utility through the joint optimization of transmission strategies and resource allocation, utilizing uplink-downlink duality, interference pricing, and a heuristic joint proportional fair scheduling algorithm. In [25], a novel BS coordination approach to mitigate inter-cell interference in orthogonal frequency-division multiple access (OFDMA) based cellular networks is proposed. By formulating the resource allocation problem as a local cooperation game among BSs, the authors demonstrate the existence of a joint-strategy Nash equilibrium that leads to global optimality of network utility. In [26], the authors investigate an energy-efficient resource allocation algorithm for multicell OFDMA systems with imperfect CSI. The proposed algorithm aims to maximize the system’s EE by jointly optimizing user scheduling, data rate adaptation, and power allocation. To address co-channel interference, multiple BSs coordinate their resource allocation strategies based on shared CSI. [27] proposes a deep reinforcement learning (RL) framework for joint power allocation and channel assignment in non-orthogonal multiple access (NOMA) systems, where two users share the same channel resources. To optimize resource allocation, the authors develop RL with an attention-based neural network (ANN) to perform channel assignment and derive a closed-form solution for power allocation. [28] investigates the joint optimization of resource allocation and user association in a heterogeneous cellular network (HCN) assisted by reconfigurable intelligent surface (RIS). The study focuses on a multi-BS multi-frequency network scenario, including 4G, 5G, millimeter wave (mmWave), and terahertz networks, with the goal of maximizing the system sum rate. To address the NP-hard and coupling relationship between RIS phase shift and resource allocation, the authors propose a block coordinate descent (BCD) method to alternately optimize local solutions and obtain the global solution. In [29], the maximization of SE in the cooperative low Earth orbit (LEO) multi-satellite network is studied, emphasizing the joint optimization of hybrid beamforming and user scheduling. In scenarios involving multiple ground users (GUs) and satellites, the proposed approach enhances SE by optimizing satellite-to-GU associations and mitigating interference using digital beamforming techniques. Different from the above works, we investigate feedback-free DL resource scheduling to facilitate multi-BS flexible cooperation in FD-RAN. Additionally, the scheduling process prioritizes meeting users’ personalized demands, incorporating priority and fairness in heavy-load network conditions, while minimizing energy costs in light-load network conditions.
@@ III-A System Model
We consider the DL network of FD-RAN illustrated in Fig. 1, where $M$ DL-BSs serve $N$ users. In FD-RAN, feedback-free MIMO transmission is employed, where the MIMO transmission parameters including the precoder, the number of spatial streams, and channel quality indicator (CQI) are determined based on the user’s geolocation through a mapping derived from historical channel data. This mapping, denoted as $\mathcal{FF}_{\text{TM}}:\bm{loc}\rightarrow TP$, relates the user’s 3-dimensional Cartesian coordinates $\bm{loc}$ to transmission parameters $TP$ according to the employed TM. TMs, such as transmit diversity, single-user MIMO (SU-MIMO), and JT, have defined various transmission methods to optimize network performance under different scenarios. In our previous work [14], we have studied how to establish such mapping, while other studies have investigated alternative approaches [15, 30]. The DL network of FD-RAN operates over a specific DL band containing $K$ orthogonal subcarriers, each of which can be used by any DL-BS to serve users. Users demanding data report their locations to the C-BS. The edge cloud collects user locations from the C-BS, makes scheduling decisions, and sends them to DL-BSs and C-BS over high-capacity, low-latency optical fibers. Then, the C-BS informs users of their assigned subcarriers via DL control information (DCI) through dedicated wireless bands, enabling the DL-BSs to transmit and the users to receive data on the specified subcarriers. DL-BSs are also informed of users’ locations, based on which they determine the appropriate precoders as well as modulation and coding schemes (MCSs) from $\mathcal{FF}_{\text{TM}}$.
@@
Fig. 2 presents the resource scheduling process over multiple time slots in the DL network of FD-RAN or 5G FDD systems. Different from FD-RAN, the process in 5G starts with the user estimating the channel and reporting CSI. In this paper, we take into account the delays associated with the reporting of location/CSI, the running time of the scheduling algorithm, and the activation of scheduling decisions, denoted as $d_{1}$, $d_{2}$, and $d_{3}$, respectively. As defined by $k_{0}$, which specifies the delay between DCI and the DL data transmission slot in the 3GPP standard [2], $d_{3}$ must be no less than the maximum delay for control signaling to reach all users such that scheduling decisions take effect simultaneously. The location/CSI reporting period and scheduling period are denoted as $T_{rp}$ and $T_{sc}$, respectively. Data transmissions depend on the previously activated scheduling decisions, while new decisions are made based on the latest received locations/CSI from the current time slot. For 5G, it can be observed that due to various delays, CSI used in actual transmission may significantly differ from that used in the most recent scheduling decision, potentially affecting the scheduling performance. Although location reporting in FD-RAN is also subject to delays, user locations generally remain stable within the ms-level scheduling period. Furthermore, FD-RAN employs fixed transmission parameters for each location, making it insensitive to transmission delays.
@@ III-A1 SU-MIMO
where $t^{\mu}=\frac{10^{-3}}{14\times 2^{\mu}}$ is the average orthogonal frequency-division multiplexing (OFDM) symbol duration in a time slot for numerology $\mu$. The subcarrier spacing in kHz is given by $2^{\mu}\times 15$. $|\mathcal{K}^{\prime}|$ denotes the number of allocated subcarriers. $OH$ represents the overhead ratio for control signaling, synchronization, and demodulation reference signals. For simplicity, we assume that selecting a higher CQI results in zero throughput. For more realistic data rates, the high-fidelity Vienna 5G link-level simulator can generate a table of BLERs for different CQI choices [35], enabling the calculation of actual rates as the original rate multiplied by $(1-\text{BLER})$.
@@ III-B Problem Formulation
Depending on the number of subcarriers allocated by multiple BSs in the cooperation set, the hybrid TM can encompass various cases, ranging from a single BS to two cooperating BSs serving the user, and extending to all BSs in the cooperation set jointly serving the user. For instance, suppose the cooperation set serving the user is $\{1,2,3\}$, and the number of subcarriers allocated by the BSs in the cooperation set is $\{8,2,4\}$. This subcarrier allocation scheme will include the following TMs: First, BSs 1, 2, and 3 jointly use 2 subcarriers for three-BS JT; then, BSs 1 and 3 use additional $4-2=2$ subcarriers to perform two-BS JT; finally, BS 1 uses the remaining $8-4=4$ subcarriers to serve the user independently. If the number of subcarriers allocated by the BSs in the cooperation set is $\{8,4,4\}$, then only the three-BS JT on 4 subcarriers and BS 1 serving the user independently with the remaining $8-4=4$ subcarriers are included in the TMs. With all decision variables determined, resource mapping is performed by allocating subcarriers to each user in turn from the first subcarrier in $\mathcal{K}$, in accordance with the aforementioned procedure.
@@ IV Resource Scheduling Under Heavy-Load Condition
Under heavy-load network conditions, the TM is fixed to JT by all DL-BSs, i.e., all DL-BSs allocate the same number of subcarriers to a given user. Thus, the decision variable for user $n$ can be simplified as $x_{n},\forall n\in\mathcal{N}$. This leads to a significant reduction in the solution space, from $(MN)^{K}$ to $N^{K}$, making the problem more tractable. To maximize the network’s weighted sum of user demand satisfaction rate while meeting the minimum demand satisfaction rate constraint, we propose a greedy algorithm to solve the problem.
@@
Given that the greedy weight of a user changes with the number of allocated subcarriers, we divide the user $n$’s information into two parts: one for the demand of the first $K_{n}^{d}-1$ subcarriers and the other for the demand of the last subcarrier. Therefore, the problem $\mathcal{Z}$ under heavy-load network conditions to be solved can be expressed as
@@
Let user $n$ be the one with the highest greedy weight. The goal is to prove that the optimal solution involves allocating as many subcarriers as possible to user $n$. We prove that this statement is true by contradiction. Suppose there exists an optimal solution in which we do not allocate as many subcarriers as possible to user $n$, and all remaining subcarriers have already been allocated. Since user $n$ has the highest greedy weight, there must exist another user $n^{\prime}$ such that $\vartheta_{n^{\prime}}<\vartheta_{n}$. We can take $\Delta$ subcarriers from the allocation to user $n^{\prime}$ and schedule them to user $n$. The change in the weighted sum of user demand satisfaction is given by $\Delta\vartheta_{n}-\Delta\vartheta_{n^{\prime}}=\Delta(\vartheta_{n}-\vartheta_{n^{\prime}})>0$, since $\vartheta_{n^{\prime}}<\vartheta_{n}$. This leads to a contradiction because the "optimal" solution assumed could be improved by reallocating some subcarriers from user $n^{\prime}$ to user $n$. Hence, it is not optimal. ∎
@@
Suppose that $\mathcal{X}$ is the optimal solution to the problem $\mathcal{Z}$ with $K_{\text{remaining}}$ subcarriers, and it has the optimal weighted sum of demand satisfaction rate $V$. Our goal is to prove that $\mathcal{X}^{\prime}=\mathcal{X}\setminus\{x_{n}^{:K_{n}^{d}-1}\}$ is the optimal solution to the subproblem $\mathcal{Z}^{\prime}=\mathcal{Z}\setminus\{\mathcal{Z}_{n}^{:K_{n}^{d}-1}\}$, which has a weighted sum of demand satisfaction rate $V^{\prime}$ and $K_{\text{remaining}}^{\prime}=K_{\text{remaining}}-(K_{n}^{d}-1)$ subcarriers available for allocation. We prove this by contradiction. Suppose that $\mathcal{X}^{\prime}$ is not the optimal solution to $\mathcal{Z}^{\prime}$ and that there exists another solution $\mathcal{X}^{\prime\prime}$ which yields a higher weighted sum of demand satisfaction rate, i.e., $V^{\prime\prime}>V^{\prime}$. Then, $\mathcal{X}^{\prime\prime}\cup\{x_{n}^{:K_{n}^{d}-1}\}$ is a feasible solution to the original problem $\mathcal{Z}$, and the corresponding weighted sum of demand satisfaction rate is $V^{\prime\prime}+v_{n}^{:K_{n}^{d}-1}>V^{\prime}+v_{n}^{:K_{n}^{d}-1}=V$. This contradicts the assumption that $V$ is optimal. ∎

### v3

# Feedback-Free Resource Scheduling for Flexible Multi-BS Cooperation in FD-RAN

@@ Abstract
Flexible cooperation among base stations (BSs) is critical to improve resource utilization efficiency and meet personalized user demands. However, its practical implementation is hindered by the current radio access network (RAN), which relies on the coupling of uplink and downlink transmissions and channel state information feedback with inherent issues such as overheads and delays. To overcome these limitations, we consider the fully-decoupled RAN (FD-RAN), in which uplink and downlink functionalities are independent, and feedback-free MIMO transmission is adopted at the physical layer. To further deliver flexible cooperation in FD-RAN, we study feedback-free downlink multi-BS resource scheduling under the practical scheduling process. The problem is considered based on network load conditions. In heavy-load states where it is impossible for all user demands to be met, an optimal greedy algorithm is proposed, maximizing the weighted sum of user demand satisfaction rates. In light-load states where at least one solution exists to satisfy all user demands, an optimal two-stage resource allocation algorithm is designed to further minimize network energy consumption by leveraging the flexibility of cooperation. Extensive simulations validate the superiority of proposed algorithms in performance and running time, and highlight the potential for realizing flexible cooperation in practice.
@@ I Introduction
However, implementing flexible cooperation confronts significant challenges since the architecture of the fifth-generation (5G) radio access network (RAN) still follows the original single-cell based design [6]. First, the coupling between uplink (UL) and downlink (DL) transmissions restricts the flexibility of BS association and resource utilization. For example, in time division duplex (TDD) systems, cooperating BSs are required to maintain consistent UL and DL frame structures, and it is not feasible to independently configure different cooperation sets for UL and DL transmissions. Moreover, the performance gains of cooperation heavily depend on the accuracy and timeliness of channel state information (CSI), which can only be ensured in systems with tightly coupled UL and DL. Second, the overheads of pilots and CSI feedback scale linearly with the number of cooperating BSs and users, reducing the effective data rate, as flexible cooperation requires CSI for an arbitrary pair of BS and user. Furthermore, the computational complexity of determining joint transmission parameters for multi-BS cooperation increases exponentially [7]. As a result, practical cooperation is generally constrained to two fixed BSs [8]. Third, inherent feedback delays in transmission lead to outdated CSI in the practical scheduling process, resulting in a mismatch between the cooperative resource scheduling decision and its desired objectives, e.g., meeting specific users’ demands.
@@
However, there still lacks existing research on feedback-free resource scheduling towards flexible multi-BS cooperation. To this end, we investigate the problem of DL cooperation set selection and subcarrier allocation in FD-RAN. The flexibility of resource allocation lies in that users can be associated with any potential DL-BSs operating in appropriate TMs. However, such flexibility also significantly increases the complexity of the problem, posing challenges to achieving optimal solutions. Furthermore, resource allocation can have diverse and even conflicting objectives, including network throughput, fairness, and energy efficiency (EE). Thus, appropriate scheduling objectives are required for various conditions to better utilize resources. Moreover, to align with the practical scheduling process, resource scheduling should meet stringent running time requirements, e.g., at the ms level. Consequently, designing resource allocation algorithms that achieve both high computational efficiency and scheduling quality in FD-RAN remains a very challenging problem.
To address the above challenges, we propose a feedback-free cooperative resource scheduling approach that can leverage the flexibility to achieve tailored scheduling objectives based on the current network load condition. We also consider resource scheduling from the user’s perspective, focusing on meeting personalized service demands. Specifically, the network conditions are classified into heavy-load and light-load based on whether all user demands can be met through JT of all DL-BSs, which provides the highest data rate per subcarrier. In heavy-load conditions, where user demands cannot be fully met, JT is still employed to maximize demand satisfaction. A greedy subcarrier allocation algorithm considering user service priorities and fairness is proposed and proves optimal. In light-load conditions, where at least one solution exists to satisfy all user demands by JT of all DL-BSs, the focus shifts to finding the most energy-efficient cooperation strategy. An efficient two-stage resource allocation (TSRA) algorithm is proposed, in which the branch-and-cut method [16] is employed to obtain the optimal solution. To improve computational efficiency, the feasibility pump algorithm [17] is incorporated to quickly generate a high-quality feasible solution as initialization. To conclude, the contributions of this paper are listed as follows:
@@
  • To solve the problems, the resource scheduler is designed with two tailored algorithms. In heavy-load conditions, a greedy subcarrier allocation algorithm is proposed and proves optimal. In light-load conditions, the TSRA algorithm is proposed to enable flexible and energy-efficient cooperation by providing a near-optimal solution within a short time.
@@ II-B Resource Allocation Towards Multi-BS Cooperation
Over the past decade, substantial efforts have been dedicated to resource allocation aimed at enabling multi-BS cooperation across diverse network settings and objectives[24, 25, 26, 27, 28, 29]. [24] proposes a coordinated scheduling, beamforming, and power allocation scheme across multiple BSs in wireless cellular networks. The study primarily aims to mitigate inter-cell interference and enhance overall network utility through the joint optimization of transmission strategies and resource allocation, utilizing uplink-downlink duality, interference pricing, and a heuristic joint proportional fair scheduling algorithm. In [25], a novel BS coordination approach to mitigate inter-cell interference in orthogonal frequency-division multiple access (OFDMA) based cellular networks is proposed. By formulating the resource allocation problem as a local cooperation game among BSs, the authors demonstrate the existence of a joint-strategy Nash equilibrium that leads to global optimality of network utility. In [26], the authors investigate an energy-efficient resource allocation algorithm for multicell OFDMA systems with imperfect CSI. The proposed algorithm aims to maximize the system’s EE by jointly optimizing user scheduling, data rate adaptation, and power allocation. To address co-channel interference, multiple BSs coordinate their resource allocation strategies based on shared CSI. [27] proposes a deep reinforcement learning (RL) framework for joint power allocation and channel assignment in non-orthogonal multiple access (NOMA) systems, where two users share the same channel resources. To optimize resource allocation, the authors develop RL with an attention-based neural network (ANN) to perform channel assignment and derive a closed-form solution for power allocation. [28] investigates the joint optimization of resource allocation and user association in a heterogeneous cellular network (HCN) assisted by reconfigurable intelligent surface (RIS). The study focuses on a multi-BS multi-frequency network scenario, including 4G, 5G, millimeter wave (mmWave), and terahertz networks, with the goal of maximizing the system sum rate. To address the NP-hard and coupling relationship between RIS phase shift and resource allocation, the authors propose a block coordinate descent (BCD) method to alternately optimize local solutions and obtain the global solution. In [29], the maximization of SE in the cooperative low Earth orbit (LEO) multi-satellite network is studied, emphasizing the joint optimization of hybrid beamforming and user scheduling. In scenarios involving multiple ground users (GUs) and satellites, the proposed approach enhances SE by optimizing satellite-to-GU associations and mitigating interference using digital beamforming techniques. Different from the above works, we investigate feedback-free DL resource scheduling to facilitate multi-BS flexible cooperation in FD-RAN. Additionally, the scheduling process prioritizes meeting users’ personalized demands and flexibly adapts to different objectives—incorporating priority and fairness under heavy-load conditions, and minimizing energy costs under light-load conditions.
@@ III-A System Model
We consider the DL network of FD-RAN illustrated in Fig. 1, where $M$ DL-BSs serve $N$ users. In FD-RAN, feedback-free MIMO transmission is employed, where transmission parameters including the precoder, the number of spatial streams, and channel quality indicator (CQI) are determined based on the user’s geolocation through a mapping derived from historical channel data. This mapping, denoted as $\mathcal{FF}_{\text{TM}}:\bm{loc}\rightarrow TP$, relates the user’s 3-dimensional Cartesian coordinates $\bm{loc}$ to transmission parameters $TP$ according to the employed TM. TMs, such as transmit diversity, single-user MIMO (SU-MIMO), and JT, have defined various transmission methods to optimize network performance under different scenarios. In our previous work [14], we have studied how to establish such mapping, while other studies have investigated alternative approaches [15, 30]. The DL network of FD-RAN operates over a specific DL band containing $K$ orthogonal subcarriers, each of which can be used by any DL-BS to serve users. Users demanding data report their locations to the C-BS. The edge cloud collects user locations from the C-BS, makes scheduling decisions, and sends them to DL-BSs and C-BS over high-capacity, low-latency optical fibers. Then, the C-BS informs users of their assigned subcarriers via DL control information (DCI) through dedicated wireless bands, enabling the DL-BSs to transmit and the users to receive data on the specified subcarriers. DL-BSs are also informed of users’ locations, based on which they determine the appropriate precoders as well as modulation and coding schemes (MCSs) from $\mathcal{FF}_{\text{TM}}$.
@@
Fig. 2 presents the resource scheduling process over multiple time slots in the DL network of FD-RAN or 5G FDD systems. Different from FD-RAN, the process in 5G starts with the user estimating the channel and reporting CSI. In this paper, we take into account the delays associated with the reporting of location/CSI, the running time of the scheduling algorithm, and the activation of scheduling decisions, denoted as $d_{1}$, $d_{2}$, and $d_{3}$, respectively. As defined by $k_{0}$, which specifies the delay between DCI and the DL data transmission slot in the 3GPP standard [2], $d_{3}$ must be no less than the maximum delay for control signaling to reach all users such that scheduling decisions can take effect simultaneously. The location/CSI reporting period and scheduling period are denoted as $T_{rp}$ and $T_{sc}$, respectively. Data transmissions depend on the previously activated scheduling decisions, while new decisions are made based on the latest received locations/CSI from the current time slot. For 5G, it can be observed that due to various delays, CSI used in actual transmission may significantly differ from that used in the most recent scheduling decision, potentially affecting the scheduling performance. Although location reporting in FD-RAN is also subject to delays, user locations generally remain stable within the ms-level scheduling period. Furthermore, FD-RAN employs fixed transmission parameters for each location, making it insensitive to transmission delays.
@@ III-A1 SU-MIMO
where $t^{\mu}=\frac{10^{-3}}{14\times 2^{\mu}}$ is the average orthogonal frequency-division multiplexing (OFDM) symbol duration in a time slot for numerology $\mu$. The subcarrier spacing in kHz is given by $2^{\mu}\times 15$. $|\mathcal{K}^{\prime}|$ denotes the number of allocated subcarriers. $OH$ represents the overhead ratio for control signaling, synchronization, and demodulation reference signals. For simplicity, we assume that selecting a higher CQI results in zero throughput. For more realistic data rates, the high-fidelity Vienna 5G link-level simulator can generate a table of BLERs for different CQI choices [35], enabling the calculation of actual rates as original rates multiplied by $(1-\text{BLER})$.
@@ III-B Problem Formulation
Depending on the number of subcarriers allocated by multiple BSs in the cooperation set, the hybrid TM can encompass various cases, ranging from a single BS to two cooperating BSs serving the user, and extending to all BSs in the cooperation set jointly serving the user. For instance, suppose the cooperation set is $\{1,2,3\}$, and the number of subcarriers allocated by the BSs in the cooperation set is $\{8,2,4\}$. This subcarrier allocation scheme will include the following TMs: First, BSs 1, 2, and 3 jointly use 2 subcarriers for three-BS JT; Then, BSs 1 and 3 use additional $4-2=2$ subcarriers to perform two-BS JT; Finally, BS 1 uses the remaining $8-4=4$ subcarriers to serve the user independently. If the number of subcarriers allocated by the BSs in the cooperation set is $\{8,4,4\}$, then only the three-BS JT on 4 subcarriers and BS 1 serving the user independently with the remaining $8-4=4$ subcarriers are included in the TMs. With all decision variables determined, resource mapping is performed by allocating subcarriers to each user in turn from the first subcarrier in $\mathcal{K}$, in accordance with the aforementioned procedure.
@@ IV Resource Scheduling Under Heavy-Load Condition
Under heavy-load network conditions, the TM is fixed to JT by all DL-BSs, i.e., all DL-BSs allocate the same number of subcarriers to a given user. Thus, the decision variable for user $n$ can be simplified as $x_{n},\forall n\in\mathcal{N}$. This leads to a significant reduction in the solution space, from $(MN)^{K}$ to $N^{K}$, making the problem more tractable. To maximize the network’s weighted sum of user demand satisfaction rates while meeting the minimum demand satisfaction rate constraint, we propose a greedy algorithm to solve the problem.
@@
Given that the greedy weight of a user changes with the number of allocated subcarriers, we divide the user $n$’s information into two parts: one for the demand of the first $K_{n}^{d}-1$ subcarriers and the other for the demand of the last subcarrier. Therefore, the problem $\mathcal{Z}$ under heavy-load network conditions can be expressed as
@@
Let user $n$ be the one with the highest greedy weight. The goal is to prove that the optimal solution involves allocating as many subcarriers as possible to user $n$. We prove that this statement is true by contradiction. Suppose there exists an optimal solution in which we do not allocate as many subcarriers as possible to user $n$, and all remaining subcarriers have already been allocated. Since user $n$ has the highest greedy weight, there must exist another user $n^{\prime}$ such that $\vartheta_{n^{\prime}}<\vartheta_{n}$. We can take $\Delta$ subcarriers from the allocation to user $n^{\prime}$ and schedule them to user $n$. The change in the weighted sum of user demand satisfaction is given by $\Delta\vartheta_{n}-\Delta\vartheta_{n^{\prime}}=\Delta(\vartheta_{n}-\vartheta_{n^{\prime}})>0$, since $\vartheta_{n^{\prime}}<\vartheta_{n}$. This leads to a contradiction because the “optimal” solution assumed could be improved by reallocating some subcarriers from user $n^{\prime}$ to user $n$. Hence, it is not optimal. ∎
@@
Suppose that $\mathcal{X}$ is the optimal solution to the problem $\mathcal{Z}$ with $K_{\text{remaining}}$ subcarriers, and it has the optimal weighted sum of demand satisfaction rates $V$. Our goal is to prove that $\mathcal{X}^{\prime}=\mathcal{X}\setminus\{x_{n}^{:K_{n}^{d}-1}\}$ is the optimal solution to the subproblem $\mathcal{Z}^{\prime}=\mathcal{Z}\setminus\{\mathcal{Z}_{n}^{:K_{n}^{d}-1}\}$, which has a weighted sum of demand satisfaction rates $V^{\prime}$ and $K_{\text{remaining}}^{\prime}=K_{\text{remaining}}-(K_{n}^{d}-1)$ subcarriers available for allocation. We prove this by contradiction. Suppose that $\mathcal{X}^{\prime}$ is not the optimal solution to $\mathcal{Z}^{\prime}$ and that there exists another solution $\mathcal{X}^{\prime\prime}$ which yields a higher weighted sum of demand satisfaction rates, i.e., $V^{\prime\prime}>V^{\prime}$. Then, $\mathcal{X}^{\prime\prime}\cup\{x_{n}^{:K_{n}^{d}-1}\}$ is a feasible solution to the original problem $\mathcal{Z}$, and the corresponding weighted sum of demand satisfaction rates is $V^{\prime\prime}+v_{n}^{:K_{n}^{d}-1}>V^{\prime}+v_{n}^{:K_{n}^{d}-1}=V$. This contradicts the assumption that $V$ is optimal. ∎